### v1

# Understanding the temperature response of biological systems: From empirical fits to mechanistic frameworks

Abstract
Virtually every biological rate changes with temperature, but the mechanisms underlying these responses differ between different processes. Here, we bring together the main theoretical approaches used to describe temperature–rate relationships, ranging from empirical curve shapes to reaction-level kinetics and network-based dynamical frameworks. These models highlight how temperature influences not only the speed of elementary reactions, but also the behavior that emerges when many reactions interact through regulation, feedback, or stochastic transitions. By outlining the assumptions and implications of each perspective, we aim to clarify how different modeling strategies connect molecular processes to physiological temperature response curves and to point toward integrative frameworks that can better explain the diversity of biological thermal responses.
preprint: APS/123-QED
Introduction
Temperature sets the pace of life by modulating molecular stability, diffusion constants, reaction rates, and material properties. From enzymatic catalysis to developmental timing, virtually every biological process depends on temperature [1, 2, 3, 4, 5]. However, our understanding of the relevant mechanisms is not complete, which makes it difficult to understand how temperature shapes biological function across scales, ranging from the motion of individual molecules to the physiology of organisms and the behavior of ecosystems (Fig. 1A) [6, 7, 8, 9, 10].
Because virtually every biologically relevant process depends on temperature, even modest temperature changes can have strong physiological effects. Many biological systems have therefore evolved mechanisms of temperature compensation, such as circadian clocks [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] and bacterial chemotaxis [23, 24], that maintain function across environmental fluctuations. Other systems exploit temperature variations to effect a desired functional response, as in fever-mediated immune responses [25] or in temperature-dependent sex determination [26, 27, 28, 29, 30, 31, 32]. Many species employing the latter strategy, such as many reptiles and some fish, are increasingly threatened by climate change [33, 34, 35, 36]. These evolutionary adaptations prove the importance of temperature effects in biological systems, but in order to understand how they work, one needs to develop a mechanistic understanding of the effect of temperature on biological systems.
The most fundamental effect of temperature on biological systems is that it determines the level of kinetic energy of the biomolecules. This is captured by the Stokes–Einstein relation $D=k_{B}T/6\pi\eta R$, which expresses the translational diffusion constant $D$ as a function of temperature $T$, viscosity $\eta$ and molecular radius $R$. The Boltzmann constant is denoted as $k_{B}$. For biomolecules with nanometer-scale dimensions in aqueous solution at body temperature, $D$ is typically on the order of $\mu m^{2}/s$. Experimental studies in E. coli have confirmed the fundamental validity of the Stokes–Einstein relation for biological systems [37]. Because biological cells work at relatively high temperatures, their diffusion constant are large and lead to frequent encounters between the biomolecules, which is essential for biochemical function and ensure that biological systems can quickly respond to environmental changes [38, 39].
Another fundamental effect of temperature is that it modulates reaction rates. Elementary chemical reactions typically follow an Arrhenius law, with reaction rates increasing exponentially as inverse temperature ($1/T$) decreases [40, 41]:
$r(T)=Ae^{-\frac{E_{a}}{RT}},$ (1)
where $E_{a}$ is the activation energy and $A$ the pre-exponential factor. The molar gas constant $R=k_{B}N_{A}$ is the molar equivalent to the Boltzmann constant and $N_{A}$ is the Avogadro constant. Strikingly, many biological processes show characteristic departures from this simple behavior: rates rise slower than expected with increasing temperatures, reach an optimum, and then decline (Fig. 1B). These deviations from simple Arrhenius scaling are commonly visualized as rate–temperature curves, which describe how a biochemical or cellular rate (e.g., an enzyme reaction, a developmental step, or a cell-cycle event) varies with inverse temperature. At the organismal and ecological levels, analogous relationships are also known as thermal performance curves (TPCs)[42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. TPCs extend the concept of rate–temperature curves to higher-level traits such as growth, fecundity, locomotor performance, behavior, or survival, quantities that emerge from the integration of many underlying physiological processes. As representative examples, in Fig. 1B we show measured rates for the growth of a M. aeruginosa cyanobacteria colony [52] and the cleavage rates during early development of D.  rerio zebrafish embryos [53]. Despite spanning different biological scales, both types of curves are typically summarized by a maximal performance or rate $r_{\mathrm{o}}$, an optimal temperature $T_{\mathrm{o}}$, a thermal breadth $W$, and the lower and upper thermal limits ($T_{\min}$, $T_{\max}$) (Fig. 1C).
Figure 1: Temperature influences biological systems across scales and can be described using different classes of models. (A) Temperature acts from the level of particles and biochemical reactions to organisms and ecosystems. (B) Examples of empirical rate–temperature relationships: growth rates of a M. aeruginosa cyanobacteria colony [52] and cleavage rates during early development of D.  rerio zebrafish embryos [53]. Both exhibit strong, nonlinear temperature dependence. (C) Conceptual illustration of Arrhenius and non-Arrhenius behavior. In an Arrhenius plot (log rate versus $1/T$), simple reactions follow a straight line, whereas biological processes typically show curvature, an optimum temperature $T_{\mathrm{o}}$, a maximal rate $r_{\mathrm{o}}$, and thermal limits ($T_{\min},T_{\max}$). (D) Overview of modeling frameworks used to describe temperature responses. Phenomenological models provide empirical fits; microscopic models derive rate–temperature relationships from reaction-level kinetics; deterministic and stochastic network models capture how temperature affects coupled biochemical or regulatory systems. All four approaches are illustrated by fitting the same zebrafish cleavage-timing dataset ([53]), demonstrating how distinct model classes can reproduce the characteristic non-Arrhenius shape of biological temperature-response curves.
A wide variety of mathematical models have been proposed to connect temperature input to systems output (Fig. 1D). Some are purely phenomenological and aim to flexibly reproduce temperature response curves. Others derive from chemical or physical principles and attempt to identify the mechanisms generating rate–temperature relationships. More recently, deterministic and stochastic network-based models have been used to explain how complex regulatory architectures respond to temperature changes. Fig. 1D applies some of these approaches to an experimental dataset of zebrafish early cleavage durations [54]. This demonstrates how phenomenological, microscopic, and network-based mechanistic approaches can each describe temperature-dependent timing in a real biological system. However, at the same time it is clear that they have different scientific content. Mechanistic models bring more predictive power, for example on the role of mutations. At the same time, they also carry the risk of missing or misrepresenting some essential molecular details [55]. In contrast, phenomenological equations or models that are independent of specific molecular processes are less likely to be falsified, but typically do not make predictions that can be tested.
Several reviews have compared empirical fitting functions and discussed their statistical performance and biological interpretation [47, 56, 46, 45]. Here, we take a different approach and contrast them to mechanistic models. We first review phenomenological models, which provide compact empirical descriptions of rate–temperature curves and TPCs and help define operational quantities such as $T_{\mathrm{o}}$, $W$, and thermal limits. We then turn to mechanistic models, organized into three levels: (i) microscopic reaction-level theories (Arrhenius, Eyring, Kramers, enzyme stability); (ii) deterministic network models, in which temperature modulates the kinetics of interconnected biochemical pathways; and (iii) stochastic network models, including Markov-chain and mean–first-passage frameworks for multistep or conformationally complex processes. Together, these perspectives demonstrate that temperature shapes biological function across scales in diverse manners, but that thermal responses typically arise from a shared basis of physical principles.
Phenomenological models
Phenomenological models provide a natural starting point for characterizing biological temperature responses. Rather than attempting to specify the underlying biochemical or biophysical mechanisms, these models focus on summarizing empirical rate–temperature relationships with flexible functional forms. Their primary aim is descriptive accuracy and parameter compression: capturing the overall shape of a rate–temperature curve or TPC using only a few biologically interpretable parameters.
Such models offer several advantages. First, they reduce complex datasets to a small set of quantities, such as $T_{\mathrm{o}}$, $r_{\mathrm{o}}$, $W$, and ($T_{\min},T_{\max}$) (Fig. 1C). This facilitates comparison across traits, species, and environments [51]. Second, phenomenological fits provide operational definitions of these quantities even when the mechanistic basis of the temperature response is unknown [57, 58]. Third, at the organismal and population levels, where traits such as growth, reproduction, and survival emerge from many underlying processes, phenomenological curves provide useful building blocks for ecological and evolutionary models that do not depend on molecular detail [59, 60, 51]. Finally, just as the Arrhenius equation historically motivated deeper theories of reaction kinetics, successful phenomenological laws can reveal robust empirical regularities that later stimulate mechanistic explanations.
[…]
Figure 2: Phenomenological and microscopic models of temperature responses. (Top) Common phenomenological approaches for fitting rate–temperature curves, including symmetric models, asymmetric models with distinct cold and warm widths, and extensions of the Arrhenius law that incorporate optimal temperatures and upper/lower limits. (Bottom) Schematic overview of key microscopic reaction–level theories: transition-state formulations such as Arrhenius and Eyring equations, Kramers’ barrier-crossing dynamics, and enzyme-catalyzed reactions with temperature-dependent active fractions.
Symmetric models
Since many empirical rate–temperature curves are unimodal, their basic shape can often be characterized by three quantities introduced before (Fig. 1B): the maximal rate $r_{\mathrm{o}}$, the optimal temperature $T_{\mathrm{o}}$, and a thermal breadth $W$. The simplest phenomenological models enforce symmetry around $T_{\mathrm{o}}$ and use these three parameters to capture the central rise and fall of the curve (Fig. 2, top left).
Gaussian. A widely used example is the Gaussian curve
$r(T)=r_{\text{o}}\exp\left(-\frac{(T-T_{\text{o}})^{2}}{2W_{\text{G}}^{2}}\right).$ (2)
which provides a good local approximation of many empirical curves near the optimum and has been used extensively as a building block in ecological and evolutionary theory [58]. Here $W_{\mathrm{G}}$ directly sets the performance breadth (width of the curve).
Quadratic polynomials. Another natural candidate to fit a rate-temperature curve with only three parameters is a quadratic polynomial. In literature this has been done for both rates and inverse rates [62, 46]. The quadratic polynomial has thermal limits that lie symmetrically around the optimal temperature while the inverse quadratic polynomial remains finite for finite temperatures and therefore does not contain thermal limits. Again, both symmetric curves seem to provide a reasonable fit of thermal data around the optimal temperature [46].
Mitchell-Angilletta model. Lastly, [46] found that a symmetric cosinuisodal symmetric temperature response curve, originally proposed by Mitchell and Angilletta as a simple mathematical model for the thermal adaptation of lizards [63], scores high on the AIC for experimental data across different traits and species, meaning that it combines a good fit with a small number of free parameters and mathematical simplicity. This curve has symmetric thermal limits, but in contrast to the quadratic polynomial, it approaches the extremal temperatures smoothly.
Together, these symmetric models provide simple three-parameter fits of the central peak and are useful when the deviations between the cold and warm sides of the curve are modest.
Asymmetric models
Empirical rate–temperature curves are, however, rarely symmetric: the decline at high temperatures is typically much sharper than at low temperatures (Fig. 2, top middle). To accommodate this asymmetry, several phenomenological families introduce separate cold and warm thermal breadths ($W_{\mathrm{c}}$ and $W_{\mathrm{h}}$).
Janisch curve. One of the earliest examples is the Janisch model (1925) [64, 65]. In first instance Janisch proposed a symmetric inverse catenary to fit the developmental rate of insects. However, because of the observed asymmetry in the data an extension with $W_{\text{c}}<W_{\text{h}}$ was developed
$r(T)=\frac{r_{0}}{2}\cdot\frac{1}{\exp\left(-\frac{T-T_{\text{o}}}{W_{\text{Jc}}}\right)+\exp\left(\frac{T-T_{\text{o}}}{W_{\text{Jh}}}\right)}.$ (3)
This curve is able to fit a variety of experimental data over a wide temperature range [66, 46].
Brière models. A more flexible and widely used family of curves is the Brière model [67], which explicitly incorporates lower and upper thermal limits. Although the original formulation was developed for the temperature dependence of arthropod development, numerous adaptations have since been introduced for other traits and taxa [67, 66, 68]. Brière-type curves consistently perform well across diverse thermal datasets—often ranking among the top AIC-scoring asymmetric models [46]. Because they naturally encode $T_{\min}$ and $T_{\max}$ while remaining parsimonious, these curves have become the primary workhorse for asymmetric phenomenological modeling.
Taylor-Sexton model. A closely related empirical form that performs well across traits and taxa is the Taylor–Sexton curve [69]. It is a fourth-order polynomial with only three free parameters $T_{\text{min}},T_{\text{o}}$ and $r_{\text{o}}$. The resulting curve approaches $T_{\text{min}}$ smoothly but not $T_{\text{max}}$. Although originally developed to model photosynthesis rates, it has since been shown to fit a broad variety of thermal datasets [46].
Matched asymptotes. More generally, Logan proposed a number of asymmetric rate–temperature curves that match different asymptotic behaviors — typically slow exponential scaling at low temperatures and a rapid exponential or polynomial decline at high temperatures — while retaining an intermediate optimal temperature [70, 71].
Phenomenological extensions of the Arrhenius law
Another route to construct a curve that fits experimental data, is to start from the observation that below the optimal temperature, performance data often scales approximately Arrhenius-like. One can then build on the Arrhenius equation to formulate a more realistic model that includes an optimal temperature and asymmetric thermal performance breadths (Fig. 2, top right).
Linearly varying $Q_{10}$. In the case of ideal Arrhenius scaling, the $Q_{10}$ factor of a process, which gives the increase in rate over a range of $10$ degrees, is constant. Atkin [72] therefore proposed to absorb deviations into a linearly temperature-dependent $Q_{10}$. This results in an asymmetric curve with interpretable limits that is easy to fit to experimental datasets [46].
Quadratic exponential. Another approach is to supplement the Arrhenius equation with a quadratic term $B$ in the exponential [73]. Formulated around a reference temperature $T^{*}$, where the rate equals $r^{*}$, this leads to
$r(T)$ $=r^{*}\exp\!\left(-E_{\mathrm{a}}\Delta\beta-B\Delta\beta^{2}\right),$ (4)
$\Delta\beta$ $=\frac{1}{RT}-\frac{1}{RT^{*}}.$ (5)
A reparametrized form with an optimum $(T_{\text{o}},r_{\text{o}})$ and a thermal breadth $W_{\text{Q}}$,
$r(T)=r_{\mathrm{o}}\exp\!\left[-\frac{T_{\mathrm{o}}^{2}}{2W_{\mathrm{Q}}^{2}}\left(\frac{T_{\mathrm{o}}}{T}-1\right)^{2}\right],$ (6)
reveals that the quadratic exponential reduces to a Gaussian with $W_{\text{G}}=W_{\text{Q}}$ near $T_{\mathrm{o}}$, linking local fits to global phenomenology.
Double exponential. A more mechanistically inspired variant is the double exponential model [74], in which biological timing is represented as the sum of two inverse Arrhenius processes with opposite‐sign activation energies $E_{a}$ and $E_{b}$
$r(T)=\frac{r_{\mathrm{o}}/2}{\exp\!\left[E_{\mathrm{a}}\!\left(\frac{1}{RT}-\frac{1}{RT_{\mathrm{o}}}\right)\right]+\exp\!\left[E_{\mathrm{b}}\!\left(\frac{1}{RT}-\frac{1}{RT_{\mathrm{o}}}\right)\right]}.$ (7)
An optimum requires $E_{\mathrm{a}}>0$ and $E_{\mathrm{b}}<0$. A local expansion around $T_{\mathrm{o}}$ yields the Janisch form (3) with $W_{\mathrm{Jc}}=R/E_{\mathrm{a}}$ and $W_{\mathrm{Jh}}=R/E_{\mathrm{b}}$, illustrating how distinct Arrhenius-based constructions collapse to similar shapes near the optimum.
Universal temperature response curve. Finally, recent work shows that any performance curve of the form $r(T)=\exp(-E_{\mathrm{a}}/RT)\,g(T)$, where $g(T)$ is a function that scales sub-exponentially below the optimal temperature, with biologically reasonable constraints can be rescaled to a universal temperature-response curve [48]:
$r(T)=r_{\mathrm{o}}\exp\!\left(\frac{T-T_{\mathrm{o}}}{W_{\mathrm{U}}}\right)\left[1-\frac{T-T_{\mathrm{o}}}{W_{\mathrm{U}}}\right],$ (8)
where $W_{\mathrm{U}}=T_{\max}-T_{\mathrm{o}}$. It suggests that many apparently distinct non-Arrhenius rate–temperature curves are mathematically related and largely characterized by the same three quantities. This is in agreement with our earlier observations for (4), (6) and (7) showing that different models lead to the same scaling around the optimal temperature, and with the work of Kontopoulos where phenomenological three-parameter curves were found to score higher on the AIC than more complex models [46].
Microscopic models
Microscopic models aim to derive rate–temperature relationships directly from the physics and chemistry of elementary reactions. They treat biological processes as thermally activated transitions, whether barrier crossing, formation of a transition state, or enzyme-mediated catalysis. In doing so they provide a mechanistic foundation for the exponential rise of rates with temperature and the deviations from Arrhenius scaling observed near thermal limits. In contrast to phenomenological fits, which summarize the shape of empirical curves, microscopic models address the origin of temperature dependence at the level of reaction steps. Below, we outline the central theoretical frameworks developed for this purpose.
Eyring theory
Roughly 40 years after Arrhenius, in 1935, Henry Eyring provided the first mechanistic foundation for the empirically observed temperature dependence of reaction rates [75]. Together with Evans and Polanyi [76], he developed transition-state theory (TST), which views reactions as equilibria between reactants and a short-lived, high-energy activated complex (Fig. 2, bottom left).
Consider $A+B\to C$. In TST, the reactants first form an unstable complex $AB^{\ddagger}$ located at the saddle point of the potential-energy surface. Formation of this complex requires free energy $\Delta G^{\ddagger}=\Delta H^{\ddagger}-T\Delta S^{\ddagger}$, with activation enthalpy $\Delta H^{\ddagger}$ and entropy $\Delta S^{\ddagger}$. Eyring’s theory yields
$r(T)=\frac{k_{B}}{h}\,e^{-\frac{\Delta G^{\ddagger}}{RT}},$ (9)
providing a physical interpretation of the Arrhenius parameters. TST captures the exponential rise of rates with temperature, yet it treats barrier crossing in thermodynamic terms. A dynamical explanation of how fluctuations drive transitions came five years later from Hendrik Kramers.
Kramers theory
Kramers analyzed thermally activated escape as a stochastic process in a potential landscape [77, 78] (Fig. 2, bottom middle). In his formulation, a particle moves under friction and random thermal kicks, described by the overdamped Langevin equation:
$\gamma\dot{x}(t)=-V^{\prime}(x)+\xi(t),$ (10)
with Gaussian noise $\xi(t)$ satisfying
$\langle\xi(t)\rangle=0,\qquad\langle\xi(t)\xi(t^{\prime})\rangle=2\gamma k_{B}T\,\delta(t-t^{\prime}).$ (11)
This stochastic trajectory corresponds to a Fokker–Planck equation with probability flux
$J(x,t):=\frac{k_{B}T}{\gamma}e^{-V/k_{B}T}\partial_{x}\!\left(e^{V/k_{B}T}p(x,t)\right)=J_{b}-J_{f},$ (12)
where $J_{f}$ gives the barrier-crossing rate. Under steady state and a saddle-point approximation one obtains
$k_{f}\simeq\frac{k_{B}T}{\gamma}\sqrt{\frac{|V^{\prime\prime}(0)|}{|V^{\prime\prime}(x_{b})|}}\exp\!\left[-\frac{V(x_{b})-V(0)}{k_{B}T}\right],$ (13)
matching the Arrhenius exponential term but with a physically defined prefactor dependent on damping and barrier curvature. Equivalent results follow from mean first-passage time calculations [79, 80]. Kramers theory thus provides a fully dynamical basis for thermally activated reaction rates.
Enzyme-catalyzed reactions
To account for nonlinear temperature scaling, microscopic treatments of enzyme-mediated processes modify classical TST by assuming (i) the reaction is controlled by a master enzyme that lowers the transition state energy barrier and (ii) the fraction of active enzyme varies with temperature (Fig. 2, bottom right). This leads to a family of modified Eyring equations with a general form:
$r(T)=P(E_{n})\,\frac{k_{B}}{h}e^{-\frac{\Delta G^{\ddagger}}{RT}},$ (14)
where $P(E_{n})$ is the fraction of enzyme in its active state. Thereby, models provided different expressions for $P(E_{n})$.
Johnson–Lewin model. Johnson and Lewin (1946) proposed that high-temperature denaturation of a master enzyme explains the downturn (“hot inactivation”) in growth-rate curves [81]. Assuming that a reversible transition between active and denatured states requires temperature-dependent free energy change $\Delta G$, they derived a double-exponential equation (instead of a single-exponential Arrhenius form), converging to $\exp(-\Delta G^{\ddagger}/(RT))$ and to $\exp(-(\Delta G^{\ddagger}-\Delta G)/(RT))$ at low and high temperatures, respectively. Following analogous reasoning, Eskil Hultin later described cold-denaturation behavior [82].
Sharpe–Schoolfield model. Sharpe and DeMichele (1977) extended this idea by allowing the enzyme to occupy one active and two inactive states—one favored at low temperatures, one at high temperatures [83]. Their expression thereby described three-phase scaling, reproducing cold and hot inactivation alongside “normal physiological range” within a single formula. Subsequent algebraic simplification produced the Sharpe–Schoolfield model [84], widely used in ecology and physiology.
Ratkowsky–Ross model. Ratkowsky, Olley, and Ross (2005) have proposed an alternative view on the thermodynamics of high- and low-temperature enzyme denaturation. Their model, motivated by experimental observations, appreciated the heat capacity of protein unfolding $\Delta C_{p}$ as a determining factor for large positive changes in free energy of protein denaturation [85, 86].
Enzyme-assisted Arrhenius (EAAR) model. The major inconsistency of the above models that can be reduced to Eq. (14), is that even in the absence of enzymes, the reaction still proceeds with the activation energy of the maximal enzyme activity, i.e., $\Delta G^{\ddagger}$. It contradicts the very purpose of the enzyme: to reduce the required energy barrier. To address this discrepancy, DeLong et al. (2016) explicitly introduced enzyme-catalyzed energy reduction into the Arrhenius equations as $E_{a}=E_{b}-E_{c}$ [87]. They expressed $E_{c}$ using the thermodynamic rules governing protein stability, similarly to the Ratkowsky-Ross model.
Deterministic network models
Mechanistic reaction–rate theories explain why particular functional forms arise at the level of single biochemical steps, but they typically treat each reaction in isolation. By contrast, biological systems rely on regulatory networks, including feedback loops, switches, and oscillatory modules. The temperature dependence of such a network cannot be inferred directly from the properties of local reactions. This motivates dynamic network-based models, in which temperature affects the parameters of a coupled system of ordinary differential equations (ODEs), and emergent properties arise from their collective nonlinear structure.
In deterministic network models, synthesis, degradation, activation, inhibition, and binding rates are all explicit temperature-dependent parameters. Even when each individual step follows a simple Arrhenius or Eyring form, the nonlinear interactions among components can produce strongly non-Arrhenius system-level behavior. These models allow one to quantify how temperature reshapes nullclines, fixed points, and bifurcation structure. For example, temperature-dependent changes in feedback strengths or reaction time scales can shift Hopf bifurcations, alter oscillation periods, suppress or induce oscillatory regimes, or generate partial temperature compensation.
Because these ODE models have a well-developed mathematical theory, they provide a natural framework for studying how networks respond dynamically to temperature changes. Bifurcation analysis, eigenvalue spectra, slow–fast decompositions, and phase–plane geometry offer mechanistic insight into how temperature alters stability, timing, and oscillatory dynamics. Such features cannot be deduced from microscopic reaction kinetics alone.
Below, we illustrate this through two representative cases: temperature scaling in the early embryonic cell cycle oscillator, and temperature compensation in chemotaxis and circadian clocks.
Temperature scaling in an embryonic cell cycle oscillator
Figure 3: Temperature response of biological oscillators. A–D: Temperature scaling in the early embryonic cell cycle. (A) Core regulatory architecture of the Xenopus embryonic cell-cycle oscillator: Cyclin B synthesis ($k_{s}$), Cdk1-dependent degradation ($k_{d}$), and a bistable activation module governing the switch-like transition between interphase and mitosis. (B) Characteristic sawtooth-like Cyclin accumulation and sharp Cdk1 activation pulses reproduced by the minimal two-ODE model of Yang et al. [88] (adapted from Rombouts et al. [54]). (C) Phase plane of the two-variable model showing Cyclin and Cdk1 nullclines, the unstable steady state (US), and the emergent relaxation-type limit cycle. (D) Temperature scaling arises from differing activation energies of synthesis ($k_{s}$) and degradation ($k_{d}$): Arrhenius plots for the two rates (left) and the corresponding nonlinear, non-Arrhenius dependence of the model-predicted oscillation period (right). E–H: Temperature compensation in a Goodwin-type circadian oscillator. (E) Minimal transcription–translation feedback architecture underlying temperature-compensated circadian rhythms [89, 20]. (F) Example time series of X and Z generated by the Goodwin oscillator, showing robust 24-h rhythms. (G) Phase portrait of the Goodwin system, illustrating the X and Z nullclines, the unstable steady state (US), and the resulting limit cycle. (H) Temperature affects the synthesis rates ($k_{X}$, $k_{Y}$, $k_{Z}$) more strongly than degradation. Because the period is governed primarily by degradation, scaling only the production terms with temperature leaves the oscillation period nearly constant.
A minimal and well-characterized example of a temperature-sensitive biochemical oscillator is the early embryonic cell cycle of the frog Xenopus laevis. Entry into mitosis (M phase) and exit back into interphase are controlled by the kinase Cdk1, which becomes active only when bound to Cyclin B and appropriately (de)phosphorylated. Active Cyclin B–Cdk1 drives mitotic events and promotes Cyclin degradation, completing a negative feedback loop.
The core regulatory architecture therefore consists of Cyclin B synthesis, its Cdk1-dependent degradation, and a bistable switch controlling Cdk1 activation (Fig. 3A). Together, the slow accumulation and degradation of Cyclin and the fast switching of Cdk1 form a classic relaxation oscillator. In its simplest formulation, this system can be described by a two-variable ODE model for Cyclin and active Cdk1 [88, 54]. The model reproduces the characteristic sawtooth-like Cyclin dynamics and sharp Cdk1 activation pulses seen in extracts (Fig. 3B). In the phase plane, the Cdk1 activation module forms an S-shaped nullcline, and its intersection with the Cyclin nullcline gives rise to an unstable state (US) around which a limit cycle emerges (Fig. 3C).
To investigate temperature scaling, Arrhenius laws can be assigned to the kinetic rates governing Cyclin synthesis ($k_{s}$), degradation ($k_{d}$), and Cdk1 activation and inactivation. Even this minimal two-ODE system is sufficient to explain non-Arrhenius period scaling. When individual reactions have different activation energies (Fig. 3D), the ratio $k_{s}/k_{d}$ changes with temperature, shifting the Cyclin nullcline in the phase plane. If $k_{s}$ increases more steeply with temperature than $k_{d}$, the Cyclin nullcline moves upward. At high enough temperatures the nullclines cease to intersect on the middle branch of the Cdk1 nullcline, the US disappears, and the limit cycle collapses into a high-Cdk1 fixed point (persistent M phase). Conversely, at low temperatures the system becomes trapped in a low-Cdk1 interphase-like state. These bifurcations generate thermal limits and a curved, non-Arrhenius period–temperature relation (Fig. 3D).
Experimental measurements in cycling Xenopus egg extracts support this mechanistic picture [54]. Within the viable embryonic temperature range, the rising (interphase) and falling (M-phase) segments of the oscillation scale differently with temperature, consistent with Cyclin synthesis being more temperature-sensitive than its degradation. This difference largely explains the non-Arrhenius scaling of the total period in the physiological range.
Temperature compensation in chemotaxis and circadian clocks
The embryonic cell cycle oscillator illustrates how differences in activation energies among reactions naturally generate non-Arrhenius scaling and thermal limits. In contrast, some biological oscillators maintain nearly constant period across a broad temperature range. Understanding this temperature compensation requires analyzing how network structure and feedback can counteract the intrinsic temperature sensitivities of individual reactions.
Temperature–robust signaling has been documented in several systems. In E. coli chemotaxis, receptors in different modified states respond oppositely to temperature changes, and the antagonistic enzymes CheR and CheB display similar temperature dependencies. Together with temperature–adjusted enzyme synthesis and degradation rates, these effects maintain a stable signaling output across the physiological range [24]. While not an oscillator, this system illustrates the general principle that compensation can arise from counterbalancing temperature influences embedded in a regulatory network.
Circadian clocks provide another striking example because here the compensatory mechanisms must not just preserve steady signaling levels but also the period of a 24h rhythm. Temperature–independent circadian periods were first noted in classic work on the fiddler crab [90] and the fruit fly [11]. One noticed that daily rhythms in locomotor activity and color changes in the crab, and daily emergence of adult flies from pupae, persisted with almost constant period despite temperature variation.
Early mechanistic proposals by Hastings and Sweeney suggested that opposing biochemical reactions within the clock could have counterbalancing temperature responses [12]. This idea motivated a series of mathematical models in which different parts of the transcription–translation feedback loop scale differently with temperature [15, 16, 18]. However, perfect compensation of temperature dependencies is not robust to parameter variation or genetic perturbations.
François, Despierre, and Siggia analyzed temperature compensation in the classic Goodwin-type transcription–translation oscillator [89, 20]. The core architecture (Fig. 3E) consists of a negative-feedback loop in which a transcription factor drives production of an mRNA X, the mRNA is translated into a protein Y, and this protein can be further converted into a second form Z that feeds back to repress transcription of X. This minimal three-variable loop generates robust 24-hour oscillations (Fig. 3F–G), and a well-defined limit cycle in phase space. In this formulation, temperature primarily affects the production (synthesis) rates of X, Y, and Z, whereas the degradation rates play the dominant role in setting the oscillation timescale. Because the period depends almost exclusively on the degradation steps, the model exhibits a form of temperature compensation whenever the degradation rates are effectively temperature independent—even if synthesis accelerates with temperature. As shown in Fig. 3H, increasing temperature strongly affects the production rates, yet the period remains close to 24 hours.
However, degradation rates are not typically temperature-independent. To address this limitation, they proposed a more robust and biologically plausible mechanism based on adaptive buffering [89, 20]. In their formulation, the molecular components of the oscillator do not exist in a single functional form but instead occupy multiple interconverting states, interpretable as distinct mRNA isoforms or post-translationally modified protein species. These interconversion reactions occur on a timescale comparable to temperature-sensitive synthesis and degradation. As temperature changes, the balance between these states shifts, and because each state contributes differently to the repression of transcription, this redistribution adjusts the effective feedback strength. In this way, the oscillator buffers the temperature dependence of the underlying biochemical rates and achieves robust temperature compensation without requiring fine-tuning of parameters. Kidd, Young, and Siggia experimentally confirmed these predictions by measuring temperature responses in Drosophila fly and mammalian systems [20]: perturbing the modification cycle disrupted compensation, whereas manipulating synthesis rates alone did not.
Recent work by Fu et al. further generalized these results, showing that such adaptive modification cycles are not model-specific but represent a generic design principle of nonlinear oscillators with period-lengthening reactions [22].
Both the embryonic cell cycle oscillator and circadian models have a similar network architecture, combining negative feedback with additional regulatory steps that change the effective feedback strength [91]. However, the consequences for temperature responses differ. In the circadian model, the interconversion between protein states occurs on timescales comparable to the temperature-sensitive synthesis and degradation reactions. This allows redistribution between isoforms to buffer temperature-induced changes and thus stabilize the period. In the embryonic cell cycle oscillator, by contrast, the dominant slow process is Cyclin accumulation, whereas switching between Cdk1 states is very fast. This strong timescale separation means that changes in the relative temperature sensitivities of synthesis and degradation directly shift the nullclines and hence the period, decreasing internal buffering. Thus, despite similarities in network structure, differences in relative timescales and in how feedback is routed lead to very different temperature-response phenotypes.
Stochastic network models
Generic temperature response of networks
Recent work has established a statistical framework for the generic temperature dependence of large biochemical networks which are stochastic and for which each rate follows an Arrhenius law [92]. Starting from the master equation of a time-continuous, space-discrete Markov process, the analysis uses the graph-theoretic representation of the mean first-passage time (MFPT) to a target state [93]. The overall rate $r$ is the inverse MFPT, and its temperature dependence $\ln{r(T)}$ can be expanded in a Taylor series whose coefficients are related to the distributions of the total activation energies along the spanning trees $E_{\mathcal{T}}$ and spanning forests of two trees $E_{\mathcal{F}}$ of the stochastic network (see Fig. 4) according to:
$\ln{r(T)}=\sum_{n=1}^{\infty}\frac{(-1)^{n}}{n!}(\kappa_{n}^{E_{\mathcal{T}}}-\kappa_{n}^{E_{\mathcal{F}}})\Delta\beta^{n}+const.,$ (15)
where $\kappa_{n}^{E_{\mathcal{T}}}$ and $\kappa_{n}^{E_{\mathcal{F}}}$ denote the $n$-th cumulants of the distributions of $E_{\mathcal{T}}$ and $E_{\mathcal{F}}$, respectively. Here, $\Delta\beta:=\frac{1}{k_{B}T}-\frac{1}{k_{B}T^{*}}$, with $T^{*}$ being a reference temperature at which the activation energies and the prefactors of the Arrhenius equations can be treated as statistically independent. Biologically, $T^{*}$ corresponds to the temperature to which the organism has evolutionarily adapted.
This result is valid for networks of any size and explains why biochemical networks often deviate from a simple linear Arrhenius relationship, as higher-order cumulants introduce systematic curvature in the Arrhenius plot. To predict the temperature response of a complex network, one has to characterize the total activation energy distributions of the spanning trees and spanning forests. Viewing the individual activation energies as independent random variables, the central limit theorem predicts that in the limit of large networks, the total activation energies approximate a normal distribution, as they are the sum of many of these individual activation energies. Then all cumulants of order $n\geq 3$ vanish, resulting in a quadratic shape in the Arrhenius plot (Fig. 4 B)
$\ln{r(T)}=(\langle E\rangle_{\mathcal{F}}-\langle E\rangle_{\mathcal{T}})\Delta\beta+\frac{\sigma_{\mathcal{T}}^{2}-\sigma_{\mathcal{F}}^{2}}{2}\Delta\beta^{2}+const,$ (16)
where $\langle E\rangle$ and $\sigma^{2}$ denote the means and variances of the respective energy distributions. This prediction agrees well with simulations of random networks and published developmental-rate data in flies which have previously been fitted by the phenomenological quadratic exponential (4) [94, 95, 73]. This limit requires that the difference in the activation energies along the trees and the forests stays remains distinct in the large network limit, typically assured by a bias in the activation energies for rates towards the target state. Moreover, it is only fully valid if the network is sufficiently complex to have a macroscopic number of spanning trees and forests.
Figure 4: Temperature response of stochastic network models. (A) Schematic representation of a generic stochastic biochemical network, where individual transitions follow Arrhenius temperature dependence. Mean first-passage times can be expressed in terms of spanning trees ($\mathcal{T}$) and spanning forests ($\mathcal{F}$), whose total activation energies determine the cumulants entering the Taylor expansion of $\ln r(T)$ [92]. (B) In large networks, the distributions of $E_{\mathcal{T}}$ and $E_{\mathcal{F}}$ become approximately Gaussian, causing all cumulants of order $n\geq 3$ to vanish and yielding a quadratic exponential dependence of the log-rate on inverse temperature. (C) A linear cascade of reversible reactions provides an analytically tractable example illustrating the emergence and breakdown of the quadratic exponential [96]. Near the reference temperature $T^{*}$, forward-biased transitions dominate and collectively generate the generic quadratic form. At extreme temperatures, however, individual critical cycles control the mean first-passage time. (D) Resulting triphasic temperature response. Around $T^{*}$, the rate follows the quadratic exponential predicted for large networks, while below $T_{+}$ and above $T_{-}$ the rate reverts to Arrhenius behavior with positive and negative effective activation energies, respectively.
A linear cascade of reversible transitions
The master equation approach has also been applied to a specific, and analytically tractable, network topology: a linear chain of $n$ reversible reactions (Fig. 4 C) [96]. Such cascades provide a reasonable representation of many multi-step biochemical processes. However, since they have only one spanning tree they do not fulfill the requirements leading to the result in (7). Still, the analysis of this model reveals that, even in this case, the generic quadratic exponential arises. Additionally, it shows how and why this quadratic scaling can break down at extreme temperatures.
Each individual step follows an Arrhenius law with a randomly distributed activation energy. At the reference temperature $T_{0}$, the cascade is assumed to be forward-biased, meaning that forward rates are, on average, larger than backward rates. Computing the MFPT from the initial to the final state shows that the overall temperature dependence of the rate falls into three regimes.
In the physiologically relevant range around $T_{0}$, the many forward transitions collectively dominate the MFPT. By the law of large numbers, the sum of their activation energies approaches a normal distribution, and the rate obeys a quadratic exponential scaling (4) with parameters
$E^{*}=\langle E\rangle,\qquad B=\frac{\sigma_{E}^{2}}{2},\qquad$ $r^{*}=\frac{n}{\langle 1/r_{f}^{*}\rangle}.$ (17)
where $E$ and $r_{f}^{*}$ denote the activation energies and forward rates of individual steps at $T^{*}$ and $\langle$$\rangle$ and $\sigma^{2}$ their mean and variance.
At low temperatures, however, the system undergoes a sudden transition. As backward reactions become comparably faster, the system spends an increasing amount of time in cycles of forward and backward reactions before reaching the final state. Below a critical temperature $T_{+}$, a single cycle dominates the MFPT. In this regime, the rate scales Arrhenius-like with a large positive activation energy that depends on the sum and difference of the activation energies within this critical cycle.
A similar breakdown of the quadratic exponential occurs at high temperatures. Above a second critical temperature $T_{-}$, one cycle again dominates the MFPT leading to an overall rate that follows the Arrhenius law. However, in this regime the effective activation energy becomes negative. The critical temperatures separating the three regimes are:
$\frac{1}{T_{\pm}}=\frac{1}{T^{*}}\pm\frac{R\sqrt{-\ln\left(\left\langle r_{f}^{*}/r_{b}^{*}\right\rangle\right)}}{\sigma_{E}},$ (18)
where $r_{f}^{*}/r_{b}^{*}$ is the ratio between individual forward and backward reaction rates at the reference temperature.
Together, these results show that the linear cascade exhibits an asymmetric triphasic temperature response: (i) a quadratic exponential near the reference temperature $T^{0}$; (ii) an Arrhenius regime with positive activation energy below $T_{-}$; and (iii) an Arrhenius regime with negative activation energy above $T_{+}$ (Fig. 4 D). This combined scaling law accurately captures the temperature dependence of a diverse set of biological processes when fitted to more than one hundred datasets that span species, traits, and timescales [96].
Discussion and conclusions
Biological processes show diverse thermal responses, yet many of their large-scale features arise from shared physical and organizational principles. The approaches reviewed here—from empirical fits to microscopic rate theory, deterministic models, and stochastic networks—illustrate that temperature not only affects individual reactions, but also the collective dynamics that result from how these reactions are connected.
Phenomenological descriptions are valuable for comparing traits and summarizing data when mechanistic detail is limited. Microscopic theories clarify why individual steps accelerate or fail under thermal stress, and why departures from simple Arrhenius behavior are expected. Network-based models, whether formulated deterministically or stochastically, reveal how feedback, branching pathways, and the statistics of multi-step processes generate characteristic nonlinearities and thermal limits at the system scale.
Together, these perspectives highlight that biological temperature responses reflect an interplay between local reaction physics and global network organization. A key direction for future work is to integrate measurements across scales, linking activation energies and molecular transitions to pathway topology and physiological performance. Such approaches will be essential for understanding past evolutionary adaptations and thermal robustness, predicting responses to fluctuating environments, assessing the sensitivity of biological systems in a warming world, and possibly even to engineer temperature-sensitive functions into biological systems.
Data and code availability
All original modeling code has been deposited at the Gelens Lab Gitlab [https://gitlab.kuleuven.be/gelenslab/publications/temperature-review], and is publicly available as of the date of publication.
Acknowledgements
The work is supported by grants from Internal funds KU Leuven (C14/23/130, LG). JBV thanks the German Academic Scholarship Foundation (Studienstiftung des Deutschen Volkes) for support.
Supplemental Material
Overview of thermal performance curves
Temperature response equations
Here, we provide an overview of the temperature response models considered in the main text and their predicted equations for rate-temperature curves.
Table 1: An overview of temperature response models considered in the main text and their predicted equations for rate-temperature curves. All parameters are explained in Table LABEL:tab:expl_model_params
Model Equation # Parameters Source
Arrhenius $r(T)={{\color[rgb]{1,.5,0}A}}\exp\!\left(-\frac{{{\color[rgb]{1,.5,0}E_{a}}}}{RT}\right)$ 22 [97]
Phenomenological models
Gaussian $r(T)={{\color[rgb]{1,.5,0}r_{\text{o}}}}\exp\!\left(-\frac{(T-{{\color[rgb]{1,.5,0}T_{\text{o}}}})^{2}}{2{{\color[rgb]{1,.5,0}W_{\text{G}}}}^{2}}\right)$ 33 [58]
Quadratic polynomial $r(T)={{\color[rgb]{1,.5,0}r_{\text{o}}}}\left(1-\frac{\left(T-{{\color[rgb]{1,.5,0}T_{\text{o}}}}\right)^{2}}{{{\color[rgb]{1,.5,0}W_{\text{Q}}}}^{2}}\right)$ 33 [62]
Inv. quadratic polynomial $r(T)={{\color[rgb]{1,.5,0}r_{\text{o}}}}\frac{{{\color[rgb]{1,.5,0}W_{\text{E}}}}^{2}}{\left(T-{{\color[rgb]{1,.5,0}T_{\text{o}}}}\right)^{2}+{{\color[rgb]{1,.5,0}W_{\text{E}}}}^{2}}$ 33 [46]
Mitchell-Angilletta $r(T)=\frac{{{\color[rgb]{1,.5,0}r_{\text{o}}}}}{2}\left(1+\cos\!\left(\frac{T-{{\color[rgb]{1,.5,0}T_{\text{o}}}}}{{{\color[rgb]{1,.5,0}W_{\text{MA}}}}}\pi\right)\right)$ 33 [63]
Janisch $r(T)=\frac{{{\color[rgb]{1,.5,0}r_{\text{o}}}}}{2}\frac{1}{\exp\!\left(-\frac{T-{{\color[rgb]{1,.5,0}T_{\text{o}}}}}{{{\color[rgb]{1,.5,0}W_{\text{Jc}}}}}\right)+\exp\!\left(\frac{T-{{\color[rgb]{1,.5,0}T_{\text{o}}}}}{{{\color[rgb]{1,.5,0}W_{\text{Jh}}}}}\right)}$ 44 [64, 65]
Brière $r(T)={{\color[rgb]{1,.5,0}a}}T(T-{{\color[rgb]{1,.5,0}T_{\text{min}}}})^{p}({{\color[rgb]{1,.5,0}T_{\text{max}}}}-T)^{q}$ 33 [67]
Simplified Brière $r(T)={{\color[rgb]{1,.5,0}a}}(T-{{\color[rgb]{1,.5,0}T_{\text{min}}}})^{p}({{\color[rgb]{1,.5,0}T_{\text{max}}}}-T)^{q}$ 33 [56, 68]
Taylor-Sexton $r(T)={{\color[rgb]{1,.5,0}r_{\text{o}}}}\left(\frac{T-{{\color[rgb]{1,.5,0}T_{\text{min}}}}}{T_{\text{o}}-{{\color[rgb]{1,.5,0}T_{\text{min}}}}}\right)^{2}\left(2-\left(\frac{T-{{\color[rgb]{1,.5,0}T_{\text{min}}}}}{{{\color[rgb]{1,.5,0}T_{\text{o}}}}-{{\color[rgb]{1,.5,0}T_{\text{min}}}}}\right)^{2}\right)$ 33 [69]
Linearly varying $Q_{10}$ $r(T)={{\color[rgb]{1,.5,0}a}}\left(\frac{{{\color[rgb]{1,.5,0}T_{\max}}}-T}{{{\color[rgb]{1,.5,0}W_{\text{Q}}}}}\right)^{T/10}$ 33 [72]
Quadratic exponential $r(T)={{\color[rgb]{1,.5,0}r^{*}}}\exp\!\left(-{{\color[rgb]{1,.5,0}E_{\text{a}}}}\left(\frac{1}{RT}-\frac{1}{RT^{*}}\right)-{{\color[rgb]{1,.5,0}B}}\left(\frac{1}{RT}-\frac{1}{RT^{*}}\right)^{2}\right)$ 33 [73]
Double exponential $r(T)=\frac{{{\color[rgb]{1,.5,0}r_{\mathrm{o}}}}/2}{\exp\!\left({{\color[rgb]{1,.5,0}E_{\mathrm{a}}^{+}}}\left(\frac{1}{RT}-\frac{1}{R{{\color[rgb]{1,.5,0}T_{\mathrm{o}}}}}\right)\right)+\exp\!\left(-{{\color[rgb]{1,.5,0}E_{\mathrm{a}}^{-}}}\left(\frac{1}{RT}-\frac{1}{R{{\color[rgb]{1,.5,0}T_{\mathrm{o}}}}}\right)\right)}.$ 4 [74]
Universal response curve $r(T)={{\color[rgb]{1,.5,0}r_{\mathrm{o}}}}\exp\!\left(\frac{T-{{\color[rgb]{1,.5,0}T_{\mathrm{o}}}}}{{{\color[rgb]{1,.5,0}W_{\mathrm{U}}}}}\right)\left(1-\frac{T-{{\color[rgb]{1,.5,0}T_{\mathrm{o}}}}}{{{\color[rgb]{1,.5,0}W_{\mathrm{U}}}}}\right)$ 3 [48]
Microscopic models
Eyring $r(T)=\frac{k_{B}}{h}\,\exp\left(\frac{{{\color[rgb]{1,.5,0}\Delta S^{\ddagger}}}}{R}-\frac{{{\color[rgb]{1,.5,0}\Delta H^{\ddagger}}}}{RT}\right)$ 22 [75]
Johnson-Lewin $r(T)=\frac{{{\color[rgb]{1,.5,0}c}}T\exp\left(-\frac{{{\color[rgb]{1,.5,0}\Delta H}}^{\ddagger}}{RT}\right)}{1+\exp\left(\frac{{{\color[rgb]{1,.5,0}\Delta S}}}{R}-\frac{{{\color[rgb]{1,.5,0}\Delta H}}}{RT}\right)}$ 44 [81]
Sharpe-DeMichele $r(T)=\frac{{{\color[rgb]{1,.5,0}c}}T\exp\left(-\frac{{{\color[rgb]{1,.5,0}\Delta H}}^{\ddagger}}{RT}\right)}{1+\exp\left(\frac{{{\color[rgb]{1,.5,0}\Delta S_{l}}}}{R}-\frac{{{\color[rgb]{1,.5,0}\Delta H_{l}}}}{RT}\right)+\exp\left(\frac{{{\color[rgb]{1,.5,0}\Delta S_{h}}}}{R}-\frac{{{\color[rgb]{1,.5,0}\Delta H_{h}}}}{RT}\right)}$ 66 [83]
Sharpe-Schoolfield $r(T)=\frac{{{\color[rgb]{1,.5,0}\rho_{25^{o}C}}}\frac{T}{298}\exp\left(\frac{{{\color[rgb]{1,.5,0}\Delta H}}^{\ddagger}}{R}\left(\frac{1}{298}-\frac{1}{T}\right)\right)}{1+\exp\left(\frac{{{\color[rgb]{1,.5,0}\Delta H_{l}}}}{R}\left(\frac{1}{{{\color[rgb]{1,.5,0}T_{1/2l}}}}-\frac{1}{T}\right)\right)+\exp\left(\frac{{{\color[rgb]{1,.5,0}\Delta H_{h}}}}{R}\left(\frac{1}{{{\color[rgb]{1,.5,0}T_{1/2h}}}}-\frac{1}{T}\right)\right)}$ 66 [84]
Ratkowsky-Ross $r(T)=\frac{{{\color[rgb]{1,.5,0}c}}T\exp\left(-\frac{{{\color[rgb]{1,.5,0}\Delta H}}^{\ddagger}}{RT}\right)}{1+\exp\left(-{{\color[rgb]{1,.5,0}n}}\frac{{{\color[rgb]{1,.5,0}\Delta H}}-T\Delta S+{{\color[rgb]{1,.5,0}\Delta C_{p}}}\!\left[(T-T_{H})-T\ln(T/T_{S})\right]}{RT}\right)}$ 55 [85, 86]
EAAR model $r(T)={{\color[rgb]{1,.5,0}A_{0}}}\exp\left(-\frac{{{\color[rgb]{1,.5,0}E_{b}}}-{{\color[rgb]{1,.5,0}E_{\Delta H}}}\left(1-\frac{T}{{{\color[rgb]{1,.5,0}T_{m}}}}\right)-{{\color[rgb]{1,.5,0}E_{\Delta Cp}}}\left(T-{{\color[rgb]{1,.5,0}T_{m}}}-T\ln\frac{T}{{{\color[rgb]{1,.5,0}T_{m}}}}\right)}{k_{B}T}\right)$ 55 [87]
Stochastic models
Generic networks $r(T)={{\color[rgb]{1,.5,0}r^{*}}}\exp\!\left(-{{\color[rgb]{1,.5,0}E^{*}}}\Delta\beta-\frac{{{\color[rgb]{1,.5,0}B}}}{2}\Delta\beta^{2}\right),$ 33 [92]
where $\Delta\beta=\frac{1}{RT}-\frac{1}{R{{\color[rgb]{0,.5,.5}T^{*}}}}$.
The macroscopic parameters can be related to the network structure:
${{\color[rgb]{1,.5,0}E^{*}}}={{\color[rgb]{0,.5,.5}\langle E\rangle_{\mathcal{T}}}}-{{\color[rgb]{0,.5,.5}\langle E\rangle_{\mathcal{F}}}}$ and ${{\color[rgb]{1,.5,0}B}}=\frac{{{\color[rgb]{0,.5,.5}\sigma_{\mathcal{T}}}}^{2}-{{\color[rgb]{0,.5,.5}\sigma_{\mathcal{F}}}}^{2}}{2}$.
Linear cascade $r(T)=\frac{1}{A^{+}\exp\!\left({{\color[rgb]{1,.5,0}E^{+}}}\Delta\beta\right)+A^{*}\exp\!\left({{\color[rgb]{1,.5,0}E^{*}}}\Delta\beta+{{\color[rgb]{1,.5,0}B}}\Delta\beta^{2}\right)+{{\color[rgb]{1,.5,0}A^{-}}}\exp\!\left({{\color[rgb]{1,.5,0}E^{-}}}\Delta\beta+{{\color[rgb]{1,.5,0}\Delta\beta_{c}}}\right)},$ 7 [96]
where $\Delta\beta=\frac{1}{RT}-\frac{1}{R{{\color[rgb]{0,.5,.5}T^{*}}}}$.
The macroscopic parameters can be related to the network structure:
${{\color[rgb]{1,.5,0}A^{*}}}={{\color[rgb]{0,.5,.5}n}}{{\color[rgb]{0,.5,.5}\langle 1/r_{f}^{*}\rangle}}$ ${{\color[rgb]{1,.5,0}E^{*}}}={{\color[rgb]{0,.5,.5}\langle E\rangle}}$, ${{\color[rgb]{1,.5,0}B}}=\frac{{{\color[rgb]{0,.5,.5}\sigma_{E}}}^{2}}{2}$
${{\color[rgb]{1,.5,0}E^{\pm}}}=\sum_{i}{{\color[rgb]{0,.5,.5}E_{fi}^{(c\pm)}}}-\sum_{i}{{\color[rgb]{0,.5,.5}E_{bi}^{(c\pm)}}}$
$\ln({{\color[rgb]{1,.5,0}A^{\pm}/{{\color[rgb]{1,.5,0}A^{*}}}}})=\pm\frac{({{\color[rgb]{0,.5,.5}E^{*}}}-{{\color[rgb]{0,.5,.5}E^{\pm}}})\sqrt{-\ln{{\color[rgb]{0,.5,.5}\langle r_{f}^{*}/{{\color[rgb]{0,.5,.5}r_{b}^{*}}}\rangle}}}}{{{\color[rgb]{0,.5,.5}\sigma_{E}}}}-\frac{\ln{{\color[rgb]{0,.5,.5}\langle r_{f}^{*}/{{\color[rgb]{0,.5,.5}r_{b}^{*}}}\rangle}}}{2}$
.0.1 Parameters of the different temperature response equations
Here, we provide an explanation of the parameters used in the different temperature response models.
Table 2: Explanation of the parameters and constants of the models in Table LABEL:Tabel_models.
Model Parameters (macroscopic and network) Constants
Arrhenius $\color[rgb]{1,.5,0}A$}}: pre-exponential factor $\text{s}^{-1}$] $R$: ideal gas constant [J $\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}E_{a}$}}: activation energy [J $\text{mol}^{-1}$]
Phenomenological models
Gaussian $\color[rgb]{1,.5,0}r_{\text{o}}$}}: optimal rate $\text{s}^{-1}$]
$\color[rgb]{1,.5,0}T_{\text{o}}$}}: optimal temperature [K]
$\color[rgb]{1,.5,0}W_{\text{G}}$}}: thermal performance breadth [K]
Quadratic polynomial $\color[rgb]{1,.5,0}r_{\text{o}}$}}: optimal rate $\text{s}^{-1}$]
$\color[rgb]{1,.5,0}T_{\text{o}}$}}: optimal temperature [K]
$\color[rgb]{1,.5,0}W_{\text{Q}}$}}: thermal performance breadth [K]
Inv. quadratic polynomial $\color[rgb]{1,.5,0}r_{\text{o}}$}}: optimal rate $\text{s}^{-1}$]
$\color[rgb]{1,.5,0}T_{\text{o}}$}}: optimal temperature [K]
$\color[rgb]{1,.5,0}W_{\text{E}}$}}: thermal performance breadth [K]
Mitchell-Angilletta $\color[rgb]{1,.5,0}r_{\text{o}}$}}: optimal rate $\text{s}^{-1}$]
$\color[rgb]{1,.5,0}T_{\text{o}}$}}: optimal temperature [K]
$\color[rgb]{1,.5,0}W_{\text{MA}}$}}: thermal performance breadth [K]
Janisch $\color[rgb]{1,.5,0}r_{\text{o}}$}}: optimal rate $\text{s}^{-1}$]
$\color[rgb]{1,.5,0}T_{\text{o}}$}}: optimal temperature [K]
$\color[rgb]{1,.5,0}W_{\text{Jh}}$}}: hot thermal performance breadth [K]
$\color[rgb]{1,.5,0}W_{\text{Jc}}$}}: cold thermal performance breadth [K]
Brière $a$: scaling factor $\text{s}^{-1}$] $p$: rising exponent (=1=1 in [67])
$\color[rgb]{1,.5,0}T_{\text{min}}$}}: upper thermal limit [K] $q$: falling exponent (=1/2=1/2 in [67])
$\color[rgb]{1,.5,0}T_{\text{max}}$}}: lower thermal limit [K]
Simplified Brière $\color[rgb]{1,.5,0}a$}}: scaling factor $\text{s}^{-1}$] $p$: rising exponent (=2=2 in [56])
$\color[rgb]{1,.5,0}T_{\text{min}}$}}: upper thermal limit [K] $q$: falling exponent (=1=1 in [56])
$\color[rgb]{1,.5,0}T_{\text{max}}$}}: lower thermal limit [K]
Taylor-Sexton $\color[rgb]{1,.5,0}r_{\text{o}}$}}: optimal rate $\text{s}^{-1}$]
$\color[rgb]{1,.5,0}T_{\text{o}}$}}: optimal temperature [K]
$\color[rgb]{1,.5,0}T_{\text{min}}$}}: lower thermal limit [K]
Linearly varying $Q_{10}$ $\color[rgb]{1,.5,0}a$}}: scaling factor $\text{s}^{-1}$]
$\color[rgb]{1,.5,0}W_{\text{Q}}$}}: thermal performance breadth [K]
$\color[rgb]{1,.5,0}T_{\text{max}}$}}: upper thermal limit [K]
Quadratic exponential $\color[rgb]{1,.5,0}r^{*}$}}: reference scaling $\text{s}^{-1}$] $R$: ideal gas constant $\text{J}\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}E^{*}$}}: activation energy [J $\text{mol}^{-1}$] $T^{*}$: reference temperature [K]
$\color[rgb]{1,.5,0}B$}}: quadratic curvature $\text{J}^{2}\text{mol}^{-2}$]
Double exponential $\color[rgb]{1,.5,0}r_{\text{o}}$}}: optimal rate $\text{s}^{-1}$] $R$: ideal gas constant $\text{J}\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}T_{\text{o}}$}}: optimal temperature [K]
$\color[rgb]{1,.5,0}E_{\text{a}}^{+}$}}: positive activation energy [J $\text{mol}^{-1}$]
$\color[rgb]{1,.5,0}E_{\text{a}}^{-}$}}: negative activation energy [J $\text{mol}^{-1}$]
Universal response curve $\color[rgb]{1,.5,0}r_{\text{o}}$}}: optimal rate $\text{s}^{-1}$]
$\color[rgb]{1,.5,0}T_{\text{o}}$}}: optimal temperature [K]
$\color[rgb]{1,.5,0}W_{\text{U}}^{+}$}}: thermal performance breadth [K]
Microscopic models
Eyring $\color[rgb]{1,.5,0}\Delta H^{\ddagger}$}}: enthalpy of activation [J mol-1] $k_{B}$: Boltzmann’s constant [J $\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}\Delta S^{\ddagger}$}}: entropy of activation [J mol-1 K-1] $h$: Planck’s constant [J $\text{mol}^{-1}\text{K}^{-1}$]
$R$: ideal gas constant [J $\text{mol}^{-1}\text{K}^{-1}$]
Johnson-Lewin $\color[rgb]{1,.5,0}\Delta H^{\ddagger}$}}: enthalpy of activation [J mol-1] $R$: ideal gas constant [J $\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}\Delta H$}}: enthalpy of denaturation [J mol-1]
$\color[rgb]{1,.5,0}\Delta S$}}: entropy of denaturation [J mol-1 K-1]
$\color[rgb]{1,.5,0}c$}}: pre-exponential factor [s-1 K-1]
Sharpe-DeMichele $\color[rgb]{1,.5,0}\Delta H^{\ddagger}$}}: enthalpy of activation [J mol-1] $R$: ideal gas constant [J $\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}\Delta H_{l}$}}: enthalpy of low-$T$ denaturation [J mol-1]
$\color[rgb]{1,.5,0}\Delta S_{l}$}}: entropy of low-$T$ denaturation [J mol-1 K-1]
$\color[rgb]{1,.5,0}\Delta H_{h}$}}: enthalpy of high-$T$ denaturation [J mol-1]
$\color[rgb]{1,.5,0}\Delta S_{h}$}}: entropy of high-$T$ denaturation [J mol-1 K-1]
$\color[rgb]{1,.5,0}c$}}: pre-exponential factor [s-1 K-1]
Sharpe-Schoolfield $\color[rgb]{1,.5,0}\Delta H^{\ddagger}$}}: enthalpy of activation [J mol-1] $R$: ideal gas constant [J $\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}\Delta H_{l}$}}: enthalpy of low-$T$ denaturation [J mol-1]
$\color[rgb]{1,.5,0}T_{1/2l}$}}: low half-denaturation temperature [K]
$\color[rgb]{1,.5,0}\Delta H_{h}$}}: enthalpy of high-$T$ denaturation [J mol-1]
$\color[rgb]{1,.5,0}T_{1/2h}$}}: high half-denaturation temperature [K]
[…]
Ratkowsky-Ross $\color[rgb]{1,.5,0}\Delta H^{\ddagger}$}}: enthalpy of activation [J mol-1] $R$: ideal gas constant [J $\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}\Delta H$}}: enthalpy of denaturation [J mol-1] $\Delta S$: entropy of denaturation [J mol-1 K-1]
$\color[rgb]{1,.5,0}\Delta C_{p}$}}: heat capacity of denaturation [J mol-1 K-1] $T_{H}$: enthalpy convergence temperature [K]
$\color[rgb]{1,.5,0}n$}}: number of enzyme’s amino acid residues $T_{S}$: entropy convergence temperature [K]
$\color[rgb]{1,.5,0}c$}}: pre-exponential factor [s-1 K-1]
EAAR model $\color[rgb]{1,.5,0}A_{0}$}}: pre-exponential factor [s-1] $k_{B}$: Boltzmann’s constant [J $\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}E_{b}$}}: baseline activation energy without catalysis [J mol-1]
$\color[rgb]{1,.5,0}E_{\Delta H}$}}: change in activation energy due to
change in denaturation enthalpy [J mol-1]
$\color[rgb]{1,.5,0}E_{\Delta Cp}$}}: change in activation energy due to
change in denaturation heat capacity [J mol-1]
$\color[rgb]{1,.5,0}T_{m}$}}: melting temperature [K]
Stochastic models
Generic networks $\color[rgb]{1,.5,0}r^{*}$}}: reference scaling $\text{s}^{-1}$] $R$: ideal gas constant $\text{J}\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}E^{*}$}}: activation energy [J $\text{mol}^{-1}$]
$\color[rgb]{1,.5,0}B$}}: quadratic curvature $\text{J}^{2}\text{mol}^{-2}$]
$\color[rgb]{0,.5,.5}T^{*}$}}: reference temperature
$\color[rgb]{0,.5,.5}\langle E\rangle_{\mathcal{T/\mathcal{F}}}$}}: average activation energy
      along trees/forests [J $\text{mol}^{-1}$]
$\color[rgb]{0,.5,.5}\sigma_{\mathcal{T/\mathcal{F}}}$}}: standard deviation activation energy
      along trees/forests [J $\text{mol}^{-1}$]
Linear cascade $\color[rgb]{1,.5,0}A^{*}$}}: reference pre-exponential factor [s] $R$: ideal gas constant $\text{J}\text{mol}^{-1}\text{K}^{-1}$]
$\color[rgb]{1,.5,0}A^{\pm}$}}: pos/neg pre-exponential factor [s]
$\color[rgb]{1,.5,0}E^{*}$}}: reference activation energy [J $\text{mol}^{-1}$]
$\color[rgb]{1,.5,0}E^{\pm}$}}: pos/ neg activation energy [J $\text{mol}^{-1}$]
$\color[rgb]{1,.5,0}B$}}: quadratic curvature $\text{J}^{2}\text{mol}^{-2}$]
$\color[rgb]{0,.5,.5}T^{*}$}}: reference temperature [K]
$\color[rgb]{0,.5,.5}n$}}: number of transitions
$\color[rgb]{0,.5,.5}r^{*}_{f}$}}: individual forward transition rate at $T^{*}$ [s-1]
$\color[rgb]{0,.5,.5}r^{*}_{b}$}}: individual backward transition rate at $T^{*}$ [s-1]
$\color[rgb]{0,.5,.5}E$}}: activation energy individual step [J $\text{mol}^{-1}$]
$\color[rgb]{0,.5,.5}E_{fi}^{c\pm}$}}: activation energy forward step
      in low/high T critical cycle [J $\text{mol}^{-1}$]
$\color[rgb]{0,.5,.5}E_{bi}^{c\pm}$}}: activation energy backward step
      in low/high T critical cycle [J $\text{mol}^{-1}$]
Mathematical details on the oscillator models
The embryonic cell cycle oscillator
The embryonic cell cycle can be described by a minimal two-variable oscillator that tracks the concentration of cyclin B (cyc) and the activity of the cyclin–Cdk1 complex (cdk1a). Following the reduced formulation of [54], cyclin is produced at a constant rate and degraded through APC/C when Cdk1 activity becomes sufficiently high, while Cdk1 activation and inactivation are governed by fast ultrasensitive feedbacks.
The dynamical equations are:
$\frac{d\,\mathrm{cyc}}{dt}$ $=k_{s}-k_{d}\,d[\mathrm{cdk1}_{a}]\;\mathrm{cyc},$ (19)
$\epsilon\,\frac{d\,\mathrm{cdk1}_{a}}{dt}$ $=k_{a}\,a[\mathrm{cdk1}_{a}]\big(\mathrm{cyc}-\mathrm{cdk1}_{a}\big)-k_{i}\,i[\mathrm{cdk1}_{a}]\,\mathrm{cdk1}_{a}.$ (20)
The ultrasensitive activation, inhibition, and APC/C–activation functions are given by:
$a[x]$ $=a_{\mathrm{Cdc25}}+b_{\mathrm{Cdc25}}\,\frac{x^{n_{\mathrm{Cdc25}}}}{K_{\mathrm{Cdc25}}^{\,n_{\mathrm{Cdc25}}}+x^{n_{\mathrm{Cdc25}}}},$ (21)
$i[x]$ $=a_{\mathrm{Wee1}}+b_{\mathrm{Wee1}}\,\frac{K_{\mathrm{Wee1}}^{\,n_{\mathrm{Wee1}}}}{K_{\mathrm{Wee1}}^{\,n_{\mathrm{Wee1}}}+x^{n_{\mathrm{Wee1}}}},$ (22)
$d[x]$ $=a_{\mathrm{APC}}+b_{\mathrm{APC}}\,\frac{x^{n_{\mathrm{APC}}}}{K_{\mathrm{APC}}^{\,n_{\mathrm{APC}}}+x^{n_{\mathrm{APC}}}}.$ (23)
For biologically realistic parameter choices, the system exhibits robust limit-cycle oscillations with a period of $\sim$30 minutes, characteristic of early embryonic cleavage cycles. In phase space, oscillations arise from slow cyclin accumulation followed by a rapid Cdk1 activation jump on the middle branch of the S-shaped Cdk1 nullcline, after which APC/C activation drives a fast relaxation phase.
Temperature dependence of the biochemical rates is introduced using Arrhenius scaling:
$k(T)=k(T_{0})\,\exp\!\left[-\frac{E_{a}}{R}\left(\frac{1}{T}-\frac{1}{T_{0}}\right)\right],$ (24)
with reference temperature
$T_{0}=18^{\circ}\mathrm{C}=291\ \mathrm{K}.$
Only the four core rate constants $k_{s}$, $k_{d}$, $k_{a}$, and $k_{i}$ are assigned nonzero activation energies; all ultrasensitive functions are treated as temperature independent. The parameters for the example shown in Fig. 3 A-E are listed in Tab. 3.
Parameter Value at $T_{0}$ $E_{a}$ (kJ mol-1)
[…]
$\epsilon$ $0.1$ 0
$a_{\mathrm{Cdc25}}$ 0.2 0
$b_{\mathrm{Cdc25}}$ 0.8 0
$K_{\mathrm{Cdc25}}$ 30nM30\ $\mathrm${nM} 0
$n_{\mathrm{Cdc25}}$ 10 0
$a_{\mathrm{Wee1}}$ 0.1 0
$b_{\mathrm{Wee1}}$ 0.4 0
$K_{\mathrm{Wee1}}$ 30nM30\ $\mathrm${nM} 0
$n_{\mathrm{Wee1}}$ 5 0
$a_{\mathrm{APC}}$ 0.1 0
$b_{\mathrm{APC}}$ 0.9 0
$K_{\mathrm{APC}}$ 30nM30\ $\mathrm${nM} 0
$n_{\mathrm{APC}}$ 15 0
Table 3: Parameters used for the embryonic cell cycle oscillator at the reference temperature $T_{0}=18^{\circ}$C. Only the core biochemical rate constants carry Arrhenius temperature dependence.
The Goodwin-type circadian oscillator
The Goodwin-type circadian oscillator shown in Fig. 3 E is governed by the following reaction equations:
$\frac{\text{d}X}{\text{d}t}$ $=k_{X}\frac{1}{1+Z^{n}}-d_{X}X,$ (25)
$\frac{\text{d}Y}{\text{d}t}$ $=k_{Y}X-d_{Y}Y,$ (26)
$\frac{\text{d}Z}{\text{d}t}$ $=k_{Z}Y-d_{Z}Z,$ (27)
where $k_{X},k_{Y},k_{Z}$ and $d_{X},d_{Y},d_{Z}$ are the synthesis and degradation rates, respectively, and $n$ is a Hill coefficient with typical values n=810n=8-10 for the Goodwin oscillator. Here, $n=9$ is used as in [89].
Introducing rescaled variables:
$x$ $=\big(\frac{k_{X}}{k_{Y}^{n}k_{Z}^{n}}\big)^{\frac{1}{n+1}}X,$ (28)
$y$ $=\big(\frac{k_{X}k_{Y}}{k_{Z}^{n}}\big)^{\frac{1}{n+1}}Y,$ (29)
$z$ $=\big(k_{X}k_{Y}k_{Z}\big)^{\frac{1}{n+1}}Z,$ (30)
in the limit $z\gg 1$, the dynamical equations read:
$\frac{\text{d}x}{\text{d}t}$ $=\frac{1}{z^{n}}-d_{X}x,$ (31)
$\frac{\text{d}y}{\text{d}t}$ $=x-d_{Y}y,$ (32)
$\frac{\text{d}z}{\text{d}t}$ $=y-d_{Z}z,$ (33)
meaning that the oscillation period only depends on the degradation rates $d_{X},d_{Y},d_{Z}$, whereas the synthesis rates $k_{X},k_{Y},k_{Z}$ only affect the amplitude. So, if the degradation rates are temperature independent, one gets an oscillation with a robust period.
As in the embryonic cell cycle model, temperature dependence is introduced through Arrhenius scaling of the synthesis rate constants:
$k=k(T=T_{0})e^{-\frac{E_{a}}{R}\big(\frac{1}{T}-\frac{1}{T_{0}}\big)},$ (34)
where $T_{0}$ is a reference temperature at which the rates are specified. The parameters for the example shown in Fig. 3 F-H are listed in Tab. 4.
rate value at $T_{0}$ $E_{a}$ [kJ mol-1]
$k_{X}$ $\kappa$ 20
$k_{Y}$ $\kappa$ 40
$k_{Z}$ $\kappa$ 60
$d_{X}$ $0.2\kappa$ 0
$d_{Y}$ $0.2\kappa$ 0
$d_{Z}$ $0.2\kappa$ 0
Table 4: The parameters used for the temperature compensated Goodwin-type model. The scale $\kappa=0.788$ $\text{h}^{-1}$ was chosen such that the period of one oscillation corresponds to a value of 2424 h. The reference temperature was set as $T_{0}=305$ K.